OpenStudy (anonymous):

Factor by grouping 12x^2-10x-18x+15

OpenStudy (anonymous):

\[12x^2-10x-18x+15=12x^2-28x+15\]

OpenStudy (anonymous):

\[12x^2−10x−18x+15\]

OpenStudy (anonymous):

\[:-10x-18x=-28x$: −10x−18x=−28x $=12x^2-28x+15$=12x2−28x+15\]

OpenStudy (mathmale):

usercode3rror: Thanks for wanting to help. Please guide the other person through these steps; do not do the work for him or her.

OpenStudy (anonymous):

okie @mathmale ill try not to do that anymore

OpenStudy (mathmale):

Jerome? Still interested in solving this problem? What factor have the first two terms in common? What factor have the last two terms in common?

OpenStudy (anonymous):

-10x and 15 12x^2 and -18x

OpenStudy (anonymous):

@mathmale